OpenStudy (jokersmart):

Help please?... What is the value of 6x12+4[(20-6)+8] ? I think 208 or 105. But I'm not sure. :\

OpenStudy (jordan123321):

hi

OpenStudy (jokersmart):

Umm.

OpenStudy (sierra!):

What do you think?

OpenStudy (sierra!):

Sorry i just saw that lol

OpenStudy (jokersmart):

it's ok

OpenStudy (sierra!):

hold on

OpenStudy (jokersmart):

Kay

OpenStudy (jdoe0001):

PEMDAS parentheses first \(\large { \begin{array}{rllll} 6\times 12+4[(20-6)+8]&\impliedby \textit{parentheses first}\\ 6\times 12+4[(14)+8]\\ 6\times 12+4[14+8]\\ 6\times 12+4[22]\\ 6\times 12+4\times 22&\impliedby \textit{multiply, from left to right}\\ 72+88&\impliedby \textit{then sum}\\ 160 \end{array} }\)

OpenStudy (sierra!):

NO DIRECT ANSWERS READ THE CODE OF CONDUCT

OpenStudy (sierra!):

YOU SHOULD KNOW THIS

OpenStudy (jokersmart):

He's showing my how to. Thanks @jdoe0001 I'll try and remember how you solved this and use it on other problems I'm at right now.

OpenStudy (jdoe0001):

hmmm ok....well... he already tried, and got 208....thus and yes, I'm aware of it, and I do take it into consideration

OpenStudy (jokersmart):

So, I'm incorrect on both answers?

OpenStudy (sierra!):

I understand he showed you how but your supposed to get the answer he is supposed to go step by step through it with you

OpenStudy (sierra!):

Yes joker

OpenStudy (jokersmart):

Ah, okay.

OpenStudy (jokersmart):

So, Simplifying 6 * 12 + 4[(20 + -6) + 8] = 0 Multiply 6 * 12 72 + 4[(20 + -6) + 8] = 0 Combine like terms: 20 + -6 = 14 72 + 4[(14) + 8] = 0 72 + 4[14 + 8] = 0 Combine like terms: 14 + 8 = 22 72 + 4[22] = 0 Multiply 4 * 22 72 + 88 = 0 Combine like terms: 72 + 88 = 160.

OpenStudy (sierra!):

really all you have to do if follow pemdas and do it left to right Parentheses Exponents Multiplication Division Addition Subtraction